# Device-Scale Atomistic Simulations of Heat Transport in Advanced Field-Effect Transistors

### v1

@@ Abstract
Self-heating in next-generation, high-power-density field-effect transistor limits performance and complicates fabrication. Here, we introduce NEP-FET, a machine-learned framework for device-scale heat transport simulations of field-effect transistors. Built upon the neuroevolution potential, the model extends a subset of the OMat24 dataset through an active-learning workflow to generate a chemically diverse, interface-rich reference set. Coupled with the FETMOD structure generator module, NEP-FET can simulate realistic field-effect transistor geometries at sub-micrometer scales containing millions of atoms, and delivers atomistic predictions of temperature fields, per-atom heat flux, and thermal stress in device structures with high fidelity. This framework enables rapid estimation of device-level metrics, including heat-flux density and effective thermal conductivity. Our results reveal pronounced differences in temperature distribution between fin-type and gate-all-around transistor architectures. The framework closes a key gap in multiscale device modeling by combining near-quantum-mechanical accuracy with device-scale throughput, providing a systematic route to explore heat transport and thermo-mechanical coupling in advanced transistors.
@@ IV Methods
The atomic energy mapping was implemented as a fully connected feedforward neural network with a single hidden layer of 80 neurons (neuron 80) and hyperbolic-tangent activation. Energies, forces and virial stresses were fitted simultaneously by minimizing a composite loss function defined as a weighted sum of the root-mean-square errors (RMSEs) of energies (per atom), forces and virials, plus $\mathcal{L}2$ regularization on the network parameters. We set equal weights for energies, forces and virials ($\lambda_{\mathrm{e}}=\lambda_{\mathrm{f}}=\lambda_{\mathrm{v}}=1$; lambda_e 1, lambda_f 1, lambda_v 1), retaining the default $\mathcal{L}2$ coefficient. Optimization used the separable natural evolution strategy built into NEP4 with a population size of 50 candidate parameter vectors per generation (population 80), and a maximum of $2\times 10^{5}$ generations (generation 200000). The final NEP-FET model was chosen as the parameter set with the lowest validation loss and attains the energy, force and stress errors reported in Fig. 2.
@@
Temperature maps were time-averaged over the steady-state window (typically 20 snapshots uniformly sampled over the last 4 ns). For visualization, the fields were lightly smoothed using a Gaussian kernel with width $\sigma=0.5$ nm; all quantitative values reported in the main text were extracted from the unsmoothed data. One-dimensional temperature profiles were obtained along annotated cutlines through the device. Interfacial temperature jumps were estimated by linear fits to bulk-like regions on either side of the interface, extrapolated to the nominal interface plane.
@@
$k_{\mathrm{eff}}=\frac{QL}{\Delta T},$ (2)
where $Q$ is the steady heat current (net thermostat power divided by the cross-sectional area normal to the transport direction), $L$ is the separation between the centroids of the hot and cold slabs, and $\Delta T$ is the fitted temperature difference between these slabs. The thermal boundary conductance of a given interface was similarly obtained as $G=J/\Delta T_{\mathrm{int}}$, where $\Delta T_{\mathrm{int}}$ is the fitted temperature jump across that interface.
Microscopic heat flux. Microscopic heat fluxes were evaluated using the following definition [Xu2025GPUMD],
$\mathbf{Q}_{i}=\sum_{j\neq i}\mathbf{r}_{ij}\left(\frac{\partial U_{j}}{\partial\mathbf{r}_{ji}}\cdot\mathbf{v}_{i}\right),$ (3)
@@
Stress fields and von Mises stress. Local stresses were obtained from the microscopic virial tensor. Assuming Newton’s third law holds, the per-atom virial is
$\bm{W}_{i}=-\frac{1}{2}\sum_{j\neq i}\mathbf{r}_{ij}\otimes\mathbf{F}_{ij},$ (4)
where $\mathbf{F}_{ij}$ is the force on atom $i$ exerted by atom $j$. The Cauchy stress tensor within a voxel of volume $V_{\text{bin}}$ was then defined as
@@ Acknowledgements.
This work was supported by the Department of Science and Technology of Liaoning Province (No. JYTMS20231613). This work also acknowledges support from the National Key R&D Project from the Ministry of Science and Technology of China (No. 2022YFA1203100), RGC GRF (No. 14220022), RGC Fund for Joint Research Labs (JLFS/E-402/24), and the CUHK PhD Studentship and Postdoctoral Fellowship.

### v2

@@ Abstract
Self-heating in next-generation, high-power-density field-effect transistor limits performance and complicates fabrication, presenting significant challenges for accurate heat transport simulations. To address this, we introduce NEP-FET, a machine-learned framework for device-scale heat transport simulations of field-effect transistors. Built upon the neuroevolution potential, the model extends a subset of the OMat24 dataset through an active-learning workflow to generate a chemically diverse, interface-rich reference set. Coupled with the FETMOD structure generator module, NEP-FET can simulate realistic field-effect transistor geometries at sub-micrometer scales containing millions of atoms, and delivers atomistic predictions of temperature fields, per-atom heat flux, and thermal stress in device structures with high fidelity. This framework enables rapid estimation of device-level metrics, including heat-flux density and effective thermal conductivity. Our results reveal pronounced differences in temperature distribution between fin-type and gate-all-around transistor architectures. The framework closes a key gap in multiscale device modeling by combining near-quantum-mechanical accuracy with device-scale throughput, providing a systematic route to explore heat transport and thermo-mechanical coupling in advanced transistors.
@@ IV Methods
The atomic energy mapping was implemented as a fully connected feedforward neural network with a single hidden layer of 80 neurons (neuron 80) and hyperbolic-tangent activation. Energies, forces and virial stresses were fitted simultaneously by minimizing a composite loss function defined as a weighted sum of the root-mean-square errors (RMSEs) of energies (per atom), forces and virials, plus $\mathcal{L}2$ regularization on the network parameters. We set equal weights for energies, forces and virials ($\lambda_{\mathrm{e}}=\lambda_{\mathrm{f}}=\lambda_{\mathrm{v}}=1$; lambda_e 1, lambda_f 1, lambda_v 1), retaining the default $\mathcal{L}2$ coefficient. Optimization used the separable natural evolution strategy built into NEP4 with a population size of 50 candidate parameter vectors per generation (population 80), and a maximum of $4\times 10^{5}$ generations (generation 200000). The final NEP-FET model was chosen as the parameter set with the lowest validation loss and attains the energy, force and stress errors reported in Fig. 2.
@@
Temperature maps were time-averaged over the steady-state window (typically 20 snapshots uniformly sampled over the last 4 ns). For visualization, the fields were lightly smoothed using a Gaussian kernel with width $\sigma=0.5$ nm; all quantitative values reported in the main text were extracted from the raw data. One-dimensional temperature profiles were obtained along annotated cutlines through the device. Interfacial temperature jumps were estimated by linear fits to bulk-like regions on either side of the interface, extrapolated to the nominal interface plane.
@@
$\kappa_{\mathrm{eff}}=\frac{QL}{\Delta T},$ (2)
where $Q$ is the steady heat flux (net thermostat power divided by the cross-sectional area normal to the transport direction), $L$ is the separation between the centroids of the hot and cold slabs, and $\Delta T$ is the fitted temperature difference between these slabs. The thermal boundary conductance of a given interface was similarly obtained as $G=J/\Delta T_{\mathrm{int}}$, where $\Delta T_{\mathrm{int}}$ is the fitted temperature jump across that interface.
Microscopic heat flux. Microscopic heat current for atom $i$ is evaluated using the following definition [Xu2025GPUMD],
$\mathbf{J}_{i}=\sum_{j\neq i}\mathbf{r}_{ij}\left(\frac{\partial U_{j}}{\partial\mathbf{r}_{ji}}\cdot\mathbf{v}_{i}\right),$ (3)
For a group of atoms within a volume $V$, the heat flux is
$\mathbf{Q}=\frac{1}{V}\sum_{i}\mathbf{J}_{i},$ (4)
@@
Stress fields and von Mises stress. Local stresses were obtained from the microscopic virial tensor. Assuming Newton’s third law holds, the per-atom virial is [Xu2025GPUMD]
$\bm{W}_{i}=\sum_{j\neq i}\mathbf{r}_{ij}\otimes\frac{\partial U_{j}}{\partial\mathbf{r}_{ji}}.$ (5)
The Cauchy stress tensor within a voxel of volume $V_{\text{bin}}$ was then defined as
@@ Acknowledgements.
This work was supported by the Department of Science and Technology of Liaoning Province (No. JYTMS20231613).